(6-sqrt(-9)-(-3-sqrt(-25)
sqrt is for the square root sign. How do you solve this equation please explain and show work

Answer by Alan3354(69443)

6-sqrt(-9)-(-3-sqrt(-25))
I added and deleted some parentheses to make them match.
----------
This can't be "solved" since it's not an equation. Equations have equal signs:
equal, equate, equation are all related.
This can be simplified.
-------
6-sqrt(-9)-(-3-sqrt(-25))
= 6 - 3i - (-3 - 5i)
= 6 - 31 + 3 + 5i
= 9 + 2i